OpenStudy (anonymous):

Solve the system of equations using the substitution method. Enter your answer as an ordered pair. 3x-y=3 y=2x+2

OpenStudy (anonymous):

You know y= 2x+2 substitute that in to the top equation. 3x-2x+2=3 subtract 2 3x-2x=1 x=1 fill that value for x into the bottom equation and solve for y. 2(1)+2=y 4=y (x,y) (1,4) :)

OpenStudy (anonymous):

thank u.

OpenStudy (anonymous):

No problem amigo.